Compute the multifactor productivity measure for an 8 hour day in which the usable output was 920 units, produced by 6 workers who used 354 pounds of materials. Workers have an hourly wage of 19 and material cost is 3.2 per pound. Overhead is 1.6 times labor cost.

The multifactor productivity (units/$) =

Blank 1

(round to 4 decimal places)
Answer #1

Output = 920 units

Each worker ges $19 per hour. So for 8 hours each worker gets $19 × 8 = $152. There are 6 workers in total. So labor cost = $152 × 6 = $912

354 pounds of material is used which cost $3.2 per pound. So material cost = $3.2 × 354 = $1132.8

Overhead costs = 1.6 times labor cost = 1.6 × $912 = $1459.2

Total cost = labor cost + material cost + overhead costs = $912 + $1132.8 + $1459.2 = $3504

Multifactor productivity = output/total cost

= 920 units / $3504

= 0.2626 units per dollar